HUBBLE SPACE TELESCOPE DAILY REPORT #991 PERIOD COVERED: 0800 EST Daily Status Report as of 285/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC/PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make one dark calibration image. The exposure lasted 1200 seconds, and the camera was pre-flashed before the observation. The observation was reported to be a normal exposure. 1.2 Completed FGS 5060 (Linking HIPPARCOS to the Extragalactic Reference Frame - Part 2 of 6) The Fine Guidance Sensor were used to make astrometric observations of the object INCA221-17. The observations will be used to determine the rotation of the reference frame used by the HIPPARCOS astrometry satellite. The astrometric observations all occurred when no telemetry was available, so the observations could not be monitored. 1.3 Completed WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make two high latitude medium-deep sky survey observations taken in parallel with the prime science observations. Both observations were received and displayed. They were reported to look fine. 1.4 Uplinked Updated Ephemeris Tables, Per MOSES 2871 The ephemeris was successfully uplinked at 284/1219Z. 1.5 Completed Two Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make a total of six internal flat calibrations images. Five images were reported to be normal, and the last had saturated corners. 1.6 Completed GHRS/2 4067 (GHRS 1ST Order Grating Spectral Lamp Test) The High Resolution Spectrograph was used to make lamp mini-function wavelength calibration observations for the first order gratings. This was strictly an internal test. No external sources were used. The observations were executed as scheduled, and no problems were reported. 1.7 Completed FGS 4893 (Orbital Motions of Close Binary Stars - cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system HD22262 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. The observation was reported to look normal. 1.8 Completed WFC 4849 (GTO High Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make two high latitude sky survey observations taken in parallel with the prime science observations. Both observations were executed as planned, and no problems were reported. 1.9 Completed Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of four observations were received. One observation was reported to be saturated, and the others were okay. 1.10 Completed FOS/RD 3791 (Quasar Absorption Line Survey: Cycle 2 Observations) The Faint object Spectrograph (Red Side) was used to make observations of quasar HS0624+690. This is a continuation of proposal FOS 2424. The first observation was received and failed. The target was just outside the aperture. A SIP corrective maneuver was successful uplinked to the spacecraft. This successfully brought the target into the aperture. Subsequent observations were reported to be okay. (See Operations Note 2.1) 1.11 Completed PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) The Planetary Camera was used to make internal flat observations to check for quantum efficiency hysteresis. Two images were taken, and both were reported to look normal. 1.12 Completed Two sets of PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) The Wide Field Camera was used to make one bias and one dark calibration images. These data are required for routine calibration of the instrument and for data reduction purposes. Both of the observations were reported to be normal calibration images. 1.13 Completed Two Sets of HSP/PMT 3373 (The Size and Composition of Planetary Ring Particles - Cycle 2 High) The High Speed Photometer was used to make observations of star 00460 as it passed behind the rings of the planet Saturn. This is a continuation of proposal 1080. The intent is to determine the size and the composition of the particles that make up the ring. The OSS monitored the observations, and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 At 284/2330Z the FGS acquisition was successful,however, the binary search target acquisition failed for FOS 3202. The science data had low counts. A corrective maneuver was successfully uplinked to the spacecraft, and subsequent science data showed much improved counts. 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 5047 (Cycle 3 FOS Discriminator Test) 3.2 Complete HSP/PMT 3373 (The Size and Composition of Planetary Ring Particles - Cycle 2 High) 3.3 Complete WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.4 Complete Three Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.5 Complete FOS/RD 3840 (The Abundance and Time Evolution of Carbon, Nitrogen, and Oxygen) 3.6 Complete WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.7 Complete PC/WFC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) 3.8 Complete Two Sets of WFC 5075 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 1) 3.9 Complete Two Sets of FOS/BL/RD 5200 (FOS Cycle 3 Scattered - Light Calibration) 3.10 Complete FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) 3.11 Complete Two Sets of PC 4714 (WF/PC 2 Calibration: PC Internal Delta Flats) 3.12 Uplink Updated Ephemeris Tables, Per MOSES 2871 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-